Jay’s is 4 times older than his son, Chris. Jay is also 36 years older than chris. How old are Jay and Chris?
Found 3 solutions by addingup, Alan3354, greenestamps:
Answer by addingup(3677)

Jay = Chris + 36
Jay = 4 x Chris
substitute for Jay:
4 x C = C + 36 = 4C = C + 36; 3C = 36; C = 12 So Chris is 12, and Jay is 4C so multiply 12 x 4 to find Jay's age

Answer by Alan3354(69443)

Jay’s is 4 times older than his son, Chris. Jay is also 36 years older than chris. How old are Jay and Chris?
------------
4 times older = 5 times as old
---
1 time older is 100% older = 2x as old.
2 times older is 200% older = 3x as old.
3 times older is 300% older = 4x as old.
4 times older is 400% older = 5x as old.
===================
J = 5C
J = C + 36
---
Sub for J
5C = C + 36
4C = 36
C = 9 yrs old
J = 36 yrs old.


Answer by greenestamps(13198)



Note the two different answers from tutors addingup and alan3354. The answer from addingup is WRONG, for the reason alan3354 says at the beginning of his response.

Unfortunately, common usage incorrectly usually interprets "Jay is 4 times OLDER THAN Chris" as meaning the same thing as "Jay is 4 times AS OLD AS Chris". But as alan3354 points out, the grammatically correct interpretation is "Jay is 5 times as old as Chris" (Chris's age, plus 4 times more Chris's age).

If this problem comes from a homework assignment, chances are the teacher thinks the wrong answer is the right answer. If so, you can show him these responses to your question to us.

The problem is widespread. I have seen problems on practice SAT exams where the incorrect interpretation is used; and just the other day I saw an example in a textbook for high school students working for the international baccalaureate degree in which the incorrect interpretation is used.